5.1 g of 1,4-di-t-butyl-2,5-dimethoxynezene (250.37 g/mol) were synthesized by reacting 9.7 mL of t-butyl alcohol (MW 74.12 g/mol, D 0.79 g/mL), 25 mL of concentrated sulfuric acid (MW 98.08 g/mol, D 1.84 g/mL), and 4.2 g of 1,4-dimethoxybenzene (MW 138.17 g/mol) together. Calculate the percent yield of this reaction. Write your answer without a percent sign and to 2 decimal places. MW=molecular weight, D=density of liquid. PLEASE SHOW WORKOH t-butanol 1,4-di-tert-butyl-2,5-dimethoxybenzene 1,4-dimethoxybenzene
Answer #1

The reaction is

               1,4-dimethyoxybenzene     +     t-butanol   ……………….> 1,4-di-tert-butyl-2,5-dimethoxybenzene

Molar mass: 138.17 g/mol                        74.12 g/mol                               250.37 g/mol

Density of t-butyl alcohol = 0.79 g/mL

Mass of t-butyl alcohol used = 9.7 x 0.79 g = 7.663 g

From the above reaction to react completely,

138.17 g of 1,4-dimethoxybenzene requires t-butanol = 74.12 g

4.2 g of 1,4-dimethoxybenzene requires t-butanol = (74.12 x 4.2)/138.17 g = 2.253 g

Hence, t-butyl alcohol is in excess and 1,4-dimethoxybenzene is the limiting reactant.

138.17 g of 1,4-dimethoxybenzene produces product = 250.37 g

4.2 g of 1,4-dimethoxybenzene produces product = (250.37 x 4.2)/138.17 g = 7.611 g

Theoretical yield = 7.611 g

Percent yield = (Actual yield/Theoretical yield) x 100% = (5.1/7.6) x 100 % = 67.11 %